# Generalization and Knowledge Transfer in Abstract Visual Reasoning Models

### v1

@@ 4 Experiments
Progressive knowledge acquisition on I-RAVEN-Mesh.
@@
In the second set of experiments we employ I-RAVEN-Mesh to examine the TL ability of each model, besides WReN that presented close-to-random performance in all previous tasks. To this end, we consider variants of partial I-RAVEN-Mesh dataset with a fraction of the training set and compare the performance of a model trained from scratch on a partial dataset to that of a model pre-trained on full I-RAVEN and fine-tuned on a part of I-RAVEN-Mesh. Fig. 6 shows that for pre-training CNN-LSTM, CoPINet, SRAN, CPCNet, and RelBase on I-RAVEN leads to gains smaller than p.p., whereas pre-training PoNG, SCL, PredRNet and STSN improved their accuracy by , , and p.p., resp. In addition, TL clearly improved performance of SCL, PredRNet and STSN in all considered settings, in particular for by , and p.p., resp., indicating their capacity for knowledge reuse.
I-RAVEN | Mesh | A/Color | A/Pos. | A/Size | A/Type | |
w/o P3 and P4 | ||||||
w/o TCN | ||||||
union |
Model | Neutral | Inter. | HO-AP | HO-TP | HO-Triples | HO-LT | HO-SC | Extra. |
PredRNet | 97.4 | 70.5 | 63.4 | 67.8 | 23.4 | 27.3 | 13.1 | 19.7 |
SCL | 87.1 | 56.0 | 79.6 | 76.6 | 23.0 | 14.1 | 12.6 | 19.8 |
PoNG (ours) | 98.1 | 75.2 | 92.1 | 97.7 | 46.1 | 16.9 | 12.6 | 19.9 |
Ablation study.
To understand the relevance of particular aspects of PoNG’s design we performed an ablation study considering several simplified model variants. The results are presented in Table 3. The removal of P3 and P4 reduces model performance, especially on A/Type ( p.p.). Similarly, disabling TCN leads to generally worse results, primarily on A/Color ( p.p.) and A/Size ( p.p.). Training without () or () typically reduces model performance, but training with one of these rule-based prediction heads compensates to some degree the lack of the other. However, the removal of both () deteriorates results across all datasets, signifying high relevance of the auxiliary training signal in PoNG’s training.
To confirm the inherent out-of-distribution generalization abilities of PoNG, we evaluated the model on all PGM regimes without performing any hyperparameter optimization (we only changed the batch size to 256 to reduce training time). Table 4 shows that PoNG achieves state-of-the-art results on PGM, particularly on the Held-out Attribute Pairs, Held-out Triple Pairs and Held-out Triples regimes, exceeding the best reference model by , and p.p., resp. We conclude that strong performance of PoNG on A-I-RAVEN and I-RAVEN-Mesh should not be attributed to any specific bias of the model towards these two datasets.
@@ 5 Conclusion
We investigate generalization capabilities of DL models in the AVR domain. To accelerate research in this area, we propose two RPM benchmarks. Attributeless-I-RAVEN introduces four generalization regimes of variable complexity that assess model’s capability to solve matrices with rules applied to novel attributes. I-RAVEN-Mesh overlays line-based patterns on top of the RPM, facilitating TL studies. Experiments on nine strong literature AVR models reveal their limitations in terms of generalization. To elevate state-of-the-art, we introduce PoNG, a novel AVR model capitalizing on parallel design, weight sharing, and normalization. PoNG outcompetes all baselines on the presented challenges and achieves significant improvement over SOTA reference models on PGM.
Acknowledgments and Disclosure of Funding
This research was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science Warsaw University of Technology. Mikołaj Małkiński was funded by the Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. This paper builds on the MSc thesis titled "Transfer learning in abstract visual reasoning domain" by Adam Kowalczyk from the Warsaw University of Technology, Warsaw, Poland.
@@ Appendix A Additional matrix examples
Appendix B Societal impact
Any effective AVR solver could be potentially misused during the online / remote IQ test assessment. Hypothetically, one might use the proposed model to unfairly claim having a higher IQ than it really is. Alternatively, the proposed datasets could be employed to develop such a model. In general, such fraudulent behaviour can lead to unequal evaluation in job recruitment or other competitive situations that relate to IQ scores.
Appendix C Limitations and future work
In this work we study generalization and knowledge transfer in AVR models employing RPM datasets. However, the set of problems in the AVR domain also includes other tasks not covered in the paper [26]. The Machine Number Sense dataset presents visual arithmetic problems [56], VAEC defines an extrapolation challenge [49], while ARC proposes a set of diverse tasks in a few-shot learning setting [6]. Future research in this area may juxtapose the performance of AVR models across a set of benchmarks oriented towards generalization to ensure generalization advances beyond the RPM datasets.
In the paper we claim that the proposed pathways block, a key component of the introduced model, is a generic module also applicable to other tasks that require reasoning over a set of objects (vector embeddings). Nevertheless, the experimental evaluation of PoNG presented in the paper is focused only on RPM benchmarks, including I-RAVEN, I-RAVEN-Mesh, Attributeless-I-RAVEN, and PGM. Assessing model’s performance on other problems constitutes an interesting extension of this work.
Appendix D Dataset details
Rule encoding.
@@ Appendix D Dataset details
As discussed in Section 3.3, we use sparse encoding [27] to represent the set of matrix rules as a vector , such that for I-RAVEN-Mesh and otherwise. The set of rules in I-RAVEN is {Constant, Progression, Arithmetic, Distribute Three} and the set of attributes is {Position, Number, Type, Size, Color}. It follows that there is unique rule–attribute pairs. In addition, the Left-Right, Up-Down, Out-InCenter, and Out-InGrid configurations in I-RAVEN comprise two components in which rules exist independently, e.g., the Left-Right component contains matrices with separate rules applied to the left and right sides. This gives an upper bound of rule–attribute combinations in each configuration. As discussed in Section 3.2 and presented in Table 1, the Mesh component introduced in I-RAVEN-Mesh comprises two attributes and four rules, leading to a total of rule–attribute combinations per configuration. As an example, in the Up-Down configuration of I-RAVEN-Mesh, there are rule–attribute combinations for the upper component, another for the lower component, and for the Mesh component. The sparse encoding encodes each rule in a matrix as a one-hot vector and applies the OR operation to the set of one-hot vectors, producing a multi-hot representation of matrix rules.
Layer | Hyperparameters |
Conv2D-ReLU-BN2D | |
Conv2D-ReLU-BN2D | |
MaxPool | |
MaxPool | |
Conv2D | |
Sum | |
Conv2D-ReLU-BN2D | |
Conv2D-ReLU-BN2D | |
MaxPool | |
MaxPool | |
Conv2D | |
Sum | |
Flatten (height & width) | |
Linear-ReLU | |
Flatten (depth & spatial) | |
LN | |
Linear-ReLU | |
LN | |
Linear | |
Sum | |
Position embedding |
Layer | Hyperparameters |
Stack | |
BN1D | |
Conv1D | |
Conv1D-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
AvgPool1D | |
BN1D | |
Conv1D | |
Conv1D-ReLU-BN1D | |
Conv1D-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
AvgPool1D | |
BN1D | |
Conv1D | |
Conv1D-ReLU-BN1D | |
Conv1D-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
G-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
GP-C-TCN-ReLU-BN1D | |
Adaptive AvgPool1D | |
Flatten (depth & feature dim) | |
Linear-ReLU-BN1D | |
Linear |
Layer | Hyperparameters |
Target head | |
Linear-ReLU-Linear | |
Linear | |
Aggregate rule head | |
Sum | |
Linear-ReLU | |
Linear | |
Target-conditioned rule head | |
Weighted sum | |
Linear |
Appendix E Model details
@@
WReN | |||
PoNG (ours) | 73.5 | 97.1 |
Test | Val | Test Val | |
CPCNet | |||
CNN-LSTM | |||
CoPINet | |||
PredRNet | |||
RelBase | |||
SCL | |||
SRAN | |||
STSN | |||
WReN | |||
PoNG (ours) | 59.4 | 96.1 |
@@
PredRNet | 98.9 | |||||||
RelBase | 98.8 | 64.3 | ||||||
SCL | ||||||||
@@
STSN | ||||||||
WReN | ||||||||
PoNG (ours) | 73.5 | 84.0 | 81.1 | 75.2 | 67.7 | 65.3 | 78.9 | 62.5 |
@@
WReN | ||||||||
PoNG (ours) | 59.4 | 58.4 | 69.7 | 65.4 | 56.6 | 56.6 | 53.2 | 55.4 |
Appendix F Extended results
Tables 13 – 13 present the results (mean and standard deviation) of all considered models on test and validation splits and the difference between these two splits. The results support the analysis of dataset difficulty presented in Section 4. The difference in model performance between test and validation splits in I-RAVEN and I-RAVEN-Mesh is negligible. In Attributeless-I-RAVEN regimes, however, the difference is significant, showing limitations of all evaluated models in terms of generalization. The biggest difference concerns the A/Type regime, suggesting that generalization of rules applied to novel shape types constitutes a real challenge for the contemporary models.
Tables 14 – 19 present the results (mean and standard deviation) of all considered models across all matrix configurations. The most challenging configurations in I-RAVEN and I-RAVEN-Mesh are 3x3Grid and Out-InGrid, in which image panels contain more objects than in the remaining configurations. Apparently such setups require stronger reasoning capabilities to correctly identify the rules applied to multiple objects. Also the results on the Left-Right and Up-Down configurations in the A/Color, A/Size, and A/Type regimes are relatively weaker. In these configurations, rules may be applied to both matrix components (left/right and up/down, resp.), which demands stronger reasoning capabilities. This also concerns the Out-InGrid configuration in the A/Position, A/Size, and A/Type regimes, and the Out-InCenter configuration in the A/Type regime. Results in the A/Position regime are close-to-perfect in configurations comprising a single object in each component (Center, Left-Right, Up-Down, and Out-InCenter) and weaker in the remaining configurations (2x2Grid, 3x3Grid and Out-InGrid). This performance drop can be attributed to the fact that Position attribute can only be effectively applied to the 2x2Grid, 3x3Grid and Out-InGrid configurations allowing modification of the object’s position. In the remaining configurations its application does not introduce any changes.
Appendix G Datasheets for datasets
In what follows, we provide the description of the introduced datasets following the Datasheets for Datasets template [11].

### v2

@@ 4 Experiments
We assess generalization of state-of-the-art models for solving RPMs on A-I-RAVEN and evaluate progressive knowledge acquisition on I-RAVEN-Mesh.
Experimental setup.
In all experiments we use the Adam optimizer Kingma and Ba (2014) with , , and a batch size set to . Learning rate is initialized to and reduced -fold (at most 3 times) if no progress is seen in the validation loss in 5 subsequent epochs, and training stops early in the case of epochs without progress. Unless stated otherwise, each model configuration was trained times with a different seed, and we report mean and standard deviation for these runs. In each experiment, we utilize training, validation, and test matrices, following the standard data split protocol taken in prior works Zhang et al. (2019a); Hu et al. (2021). All models are trained with the auxiliary loss with sparse encoding Małkiński and Mańdziuk (2024a) and . Experiments were run on a worker with a single NVIDIA DGX A100 GPU.
Models.
In addition to the simple CNN-LSTM baseline Barrett et al. (2018), we assess generalization of SOTA AVR models including WReN Barrett et al. (2018), CoPINet Zhang et al. (2019b), RelBase Spratley et al. (2020), SCL Wu et al. (2020), MRNet Benny et al. (2021), ALANS Zhang et al. (2021), SRAN Hu et al. (2021), PrAE Zhang et al. (2022), CPCNet Yang et al. (2023b), PredRNet Yang et al. (2023a), STSN Mondal et al. (2023), and DRNet Zhao et al. (2024). For direct comparison, we evaluate all models on I-RAVEN following the above-described experimental setup.
Reproducibility.
To guarantee reproducibility of experiments, we use a fixed set of random seeds and turn off hardware and framework features concerning indeterministic computation wherever possible. Together with the code, we provide the full training script that can be used to run all training jobs. The training job is packaged as a Docker image with fixed dependencies to isolate the configuration of the training environment. The released code allows for generation of all datasets from scratch, eliminating the dependency on file-hosting services required to distribute the data. The code for reproducing all experiments is publicly accessible at: https://github.com/mikomel/raven
4.1 Generalization on A-I-RAVEN
Main regimes.
In the first set of experiments we evaluate all considered models on primary generalization regimes of A-I-RAVEN. The results are presented in Table 2, along with the reference results on I-RAVEN and I-RAVEN-Mesh. The best outcomes on A/Color and A/Position are achieved by DRNet, followed by RelBase and SCL that perform comparably. In A/Size and A/Type, SCL outperforms other models, with DRNet and RelBase taking the second and third place, resp. Interestingly, the top 3 models present a mix of architectures comprising large models, such as DRNet that includes a Vision Transformer backbone (M params), as well as small models, such as SCL and RelBase that include mainly convolutional and feed-forward layers (M and M params, resp.). This suggests that various architectural approaches may be taken to achieve reasonable generalization performance in solving RPMs.
Extended regimes.
Table 3 shows the aggregated performance of all considered models on extended A-I-RAVEN regimes. Similarly to the main regimes, the best results are achieved by SCL, RelBase, and DRNet. Overall, replacing the Constant rule in the training set of the A/Color regime with Progression yields a dataset of slightly lower complexity, as the best model on A/Color-Progression achieved accuracy, a p.p. increase compared to the best result on A/Color. Conversely, using the Arithmetic and Distribute Three rules increases the difficulty, as measured by the drop of the max accuracy by p.p. and p.p., resp. Furthermore, using a pair of held-out attributes significantly increases the complexity. For instance, in A/ColorType, the most challenging regime, the best result is only . We conclude that A-I-RAVEN provides a suite of challenging regimes of variable complexity, in which even the best-performing models are far from solving all test matrices.
Dataset difficulty.
Across all A-I-RAVEN regimes, the highest average result was achieved by SCL (), followed by DRNet () and RelBase (). While SCL achieved test accuracy on I-RAVEN, on A-I-RAVEN regimes it scored from on A/ColorType to on A/Position. Similar differences can be observed for all remaining models, which shows that generalization regimes of A-I-RAVEN pose a bigger challenge than the base dataset.
Fig. 4 displays the difference in performance of top-3 models on test and validation splits. On I-RAVEN and I-RAVEN-Mesh the difference is negligible, as in these datasets both splits follow the same distribution. However, the difference in attributeless regimes is significant, indicating the need for further research on generalization.
Tables 4 – 15 in Appendix C present the results of all considered models on test and validation splits and the difference between these two splits for particular datasets/regimes. The difference in model performance between test and validation splits in I-RAVEN (Table 4) and I-RAVEN-Mesh (Table 5) is negligible. In A-I-RAVEN regimes, however, the difference is significant, showing limitations of all evaluated models in terms of generalization. Across primary regimes (Tables 6 – 9), the biggest difference concerns the A/Type regime, suggesting that generalization of rules applied to novel shape types constitutes a real challenge for the contemporary models. In all extended regimes concerning held-out attribute pairs (A/ColorSize, A/ColorType, and A/SizeType) the performance difference on test and validation splits is bigger than in the primary regimes (see Tables 10 – 12). This drop stems from overall weaker performance on the test split, confirming high difficulty of these regimes. Model performance on the next regimes concerning the Color attribute and rules other than Constant (A/Color-Progression, A/Color-Arithmetic, and A/Color-DistributeThree) is better, though further progress in generalization is required to fully close the performance gap between test and validation splits (see Tables 13 – 15).
Per-configuration results.
Tables 16 – 27 in Appendix C present the detailed results of all considered models for all matrix configurations. The most challenging configurations in I-RAVEN and I-RAVEN-Mesh are 3x3Grid and Out-InGrid, in which image panels contain more objects than in the remaining configurations. Apparently, such setups require stronger reasoning capabilities to correctly identify the rules applied to multiple objects. Also, the results on the Left-Right and Up-Down configurations are relatively weaker in most regimes. In these configurations, rules may be applied to both matrix components (left/right and up/down, resp.), increasing the task complexity. This also concerns the Out-InGrid configuration in the A/Size regime, and the Out-InCenter configuration in the A/SizeType regime. Results in the A/Position regime are close-to-perfect in configurations comprising a single object in each component (Center, Left-Right, Up-Down, and Out-InCenter) and weaker in the remaining configurations (2x2Grid, 3x3Grid and Out-InGrid). This performance drop can be attributed to the fact that Position attribute can only be effectively applied to the 2x2Grid, 3x3Grid and Out-InGrid configurations allowing modification of the object’s position. In the remaining configurations its application does not introduce any changes.
4.2 Progressive knowledge acquisition on I-RAVEN-Mesh
@@
In the second set of experiments we employ I-RAVEN-Mesh to examine the TL ability of the best performing models. To this end, we consider variants of partial I-RAVEN-Mesh dataset with a fraction of the training set and compare the performance of a model trained from scratch on a partial dataset to that of a model pre-trained on full I-RAVEN and fine-tuned on the respective part of I-RAVEN-Mesh. Fig. 5 shows that for pre-training RelBase, MRNet, CPCNet, SRAN, and CoPINet on I-RAVEN leads to gains smaller than p.p., whereas pre-training DRNet, SCL, PredRNet, and STSN improved their accuracy by , , and p.p., resp. In addition, TL clearly improved performance of DRNet, SCL, PredRNet, and STSN in all considered settings. In particular for by , , , and p.p., resp., indicating the models’ capacity for knowledge reuse.
@@ 5 Conclusion
We investigate generalization capabilities of DL models in the AVR domain. To accelerate research in this area, we propose two RPM benchmarks. A-I-RAVEN introduces generalization regimes of variable complexity that assess model’s capability to solve matrices with rules applied to novel attributes at various levels of complexity (primary and extended regimes). Contrary to the existing PGM dataset, A-I-RAVEN features compositionality, offers a variety of figure configurations, and above all does not require substantial computational resources. I-RAVEN-Mesh overlays line-based patterns on top of the RPM, facilitating TL studies. Experiments on strong literature AVR models reveal their limitations in terms of generalization. We believe that the introduced datasets complement existing RPM benchmarks and will foster progress in the AVR area.
Limitations and future work.
In this work we study generalization and knowledge transfer in contemporary AVR models employing RPM datasets. While RPMs are by far the most popular AVR tasks, the AVR domain also includes other types of problems not covered in the paper Małkiński and Mańdziuk (2023). The Machine Number Sense dataset presents visual arithmetic problems Zhang et al. (2020), VAEC defines an extrapolation challenge Webb et al. (2020), while ARC proposes a set of diverse tasks in a few-shot learning setting Chollet (2019). Similar studies could be performed on problems other than RPMs to test the performance and knowledge transfer abilities of AVR models in other problem settings.
Acknowledgments
This research was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science Warsaw University of Technology. Mikołaj Małkiński was funded by the Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. This paper builds on the MSc thesis titled ”Transfer learning in abstract visual reasoning domain” by Adam Kowalczyk from the Warsaw University of Technology, Warsaw, Poland.
@@ Appendix A Additional matrix examples
Figure 6 presents matrix examples from A/Type and A/Size, the primary regimes of A-I-RAVEN. Figures 7, 8 and 9 depict matrix examples from the extended regimes of A-I-RAVEN: A/ColorSize and A/ColorType (Fig. 7), A/SizeType and A/Color-Progression (Fig. 8), and A/Color-Arithmetic and A/Color-DistributeThree (Fig. 9). Figure 10 presents matrix examples from I-RAVEN-Mesh concerning the Number attribute.
Appendix B Auxiliary training
@@
As discussed in Section 4, we use sparse encoding Małkiński and Mańdziuk [2024a] to represent the set of matrix rules as a vector , such that for I-RAVEN-Mesh and otherwise. The set of rules in I-RAVEN is {Constant, Progression, Arithmetic, Distribute Three} and the set of attributes is {Position, Number, Type, Size, Color}. It follows that there is unique rule–attribute pairs. In addition, the Left-Right, Up-Down, Out-InCenter, and Out-InGrid configurations in I-RAVEN comprise two components in which rules exist independently, e.g., the Left-Right component contains matrices with separate rules applied to the left and right sides. This gives an upper bound of rule–attribute combinations in each configuration.
As discussed in Section 3.2 and presented in Table 1, the Mesh component introduced in I-RAVEN-Mesh comprises two attributes and four rules, leading to a total of rule–attribute combinations per configuration. As an example, in the Up-Down configuration of I-RAVEN-Mesh, there are rule–attribute combinations for the upper component, another for the lower component, and for the Mesh component.
The sparse encoding encodes each rule in a matrix as a one-hot vector and applies the OR operation to the set of one-hot vectors, producing a multi-hot representation of matrix rules.
Appendix C Extended results
Tables 4 – 15 present the results (mean and standard deviation) of all considered models on test and validation splits and the difference between these two splits for particular datasets/regimes.
Tables 16 – 27 present the results (mean and standard deviation) of all considered models in detail for all matrix configurations.
@@ Appendix C Extended results
| DRNet | |||
| MRNet | |||
| PrAE | |||
| PredRNet | |||
| RelBase | 84.9 | 85.0 | |
| SCL | |||
| SRAN | |||
| STSN | |||
| WReN |
| Test | Val | Test Val | |
| ALANS | |||
| CPCNet | |||
| CNN-LSTM | |||
| CoPINet | |||
| DRNet | 70.0 | 95.4 | |
| MRNet | |||
| PrAE | |||
@@
| DRNet | |||
| MRNet | |||
| PrAE | |||
| PredRNet | |||
| RelBase | |||
| SCL | 65.6 | 94.0 | |
| SRAN | |||
| STSN | |||
| WReN |
| Test | Val | Test Val | |
| ALANS | |||
| CPCNet | |||
| CNN-LSTM | |||
| CoPINet | |||
| DRNet | |||
| MRNet | |||
| PrAE | |||
| PredRNet | |||
| RelBase | |||
| SCL | 49.5 | 95.9 | |
| SRAN | |||
| STSN | |||
| WReN |
| Test | Val | Test Val | |
| ALANS | |||
| CPCNet | |||
| CNN-LSTM | |||
| CoPINet | |||
| DRNet | 95.9 | ||
| MRNet | |||
| PrAE | |||
| PredRNet | |||
| RelBase | |||
| SCL | 40.8 | ||
| SRAN | |||
| STSN | |||
| WReN |
| Test | Val | Test Val | |
| ALANS | |||
| CPCNet | |||
| CNN-LSTM | |||
| CoPINet | |||
| DRNet | |||
| MRNet | |||
| PrAE | |||
| PredRNet | |||
| RelBase | |||
| SCL | 32.0 | 96.4 | |
| SRAN | |||
| STSN | |||
| WReN |
| Test | Val | Test Val | |
| ALANS | |||
| CPCNet | |||
| CNN-LSTM | |||
| CoPINet | |||
| DRNet | |||
| MRNet | 97.7 | ||
| PrAE | |||
| PredRNet | |||
| RelBase | |||
| SCL | 33.5 | ||
| SRAN | |||
| STSN | |||
| WReN |
| Test | Val | Test Val | |
| ALANS | |||
| CPCNet | |||
| CNN-LSTM | |||
| CoPINet | |||
| DRNet | 96.1 | ||
| MRNet | |||
| PrAE | |||
| PredRNet | |||
| RelBase | |||
| SCL | 75.6 | ||
| SRAN | |||
| STSN | |||
| WReN |
| Test | Val | Test Val | |
| ALANS | |||
| CPCNet | |||
| CNN-LSTM | |||
| CoPINet | |||
| DRNet | 66.7 | 93.5 | |
| MRNet | |||
| PrAE | |||
@@
| WReN |
@@
| RelBase | 66.8 | |||||||
| SCL | 71.3 | 73.5 | ||||||
| SRAN | ||||||||
| STSN | ||||||||
| WReN |
| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
| ALANS | ||||||||
| CPCNet | ||||||||
| CNN-LSTM | ||||||||
| CoPINet | ||||||||
| DRNet | 77.5 | 99.4 | 99.2 | 65.4 | ||||
| MRNet | ||||||||
| PrAE | 56.5 | 50.2 | ||||||
| PredRNet | 98.7 | 98.9 | ||||||
@@
| DRNet | ||||||||
| MRNet | ||||||||
| PrAE | ||||||||
| PredRNet | ||||||||
| RelBase | ||||||||
| SCL | 65.6 | 66.2 | 72.1 | 67.0 | 61.7 | 62.0 | 68.6 | 61.5 |
@@
| STSN | 48.5 | |||||||
| WReN |
@@
| DRNet | 33.0 | |||||||
| MRNet | ||||||||
| PrAE | ||||||||
| PredRNet | ||||||||
| RelBase | 37.1 | |||||||
| SCL | 40.8 | 40.8 | 49.2 | 45.7 | 43.4 | 46.0 | ||
| SRAN | ||||||||
| STSN | ||||||||
| WReN |
| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
| ALANS | ||||||||
| CPCNet | ||||||||
| CNN-LSTM | ||||||||
| CoPINet | ||||||||
| DRNet | 38.0 | |||||||
| MRNet | ||||||||
| PrAE | 30.3 | 28.1 | ||||||
| PredRNet | ||||||||
| RelBase | 27.5 | |||||||
| SCL | 32.0 | 25.9 | 44.4 | 40.1 | ||||
| SRAN | ||||||||
| STSN | ||||||||
| WReN |
| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
| ALANS | ||||||||
| CPCNet | ||||||||
| CNN-LSTM | ||||||||
| CoPINet | ||||||||
| DRNet | 29.4 | |||||||
| MRNet | ||||||||
| PrAE | ||||||||
@@
| WReN |
| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
| ALANS | ||||||||
| CPCNet | ||||||||
| CNN-LSTM | ||||||||
| CoPINet | ||||||||
| DRNet | 85.4 | 77.6 | ||||||
| MRNet | ||||||||
| PrAE | 75.0 | 83.0 | 82.9 | |||||
| PredRNet | ||||||||
| RelBase | ||||||||
| SCL | 75.6 | 84.0 | 66.4 | |||||
| SRAN | ||||||||
| STSN | ||||||||
| WReN |
| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
| ALANS | ||||||||
| CPCNet | ||||||||
| CNN-LSTM | ||||||||
| CoPINet | ||||||||
| DRNet | 66.7 | 62.5 | 63.3 | 63.2 | 77.5 | |||
| MRNet | ||||||||
| PrAE | ||||||||
| PredRNet | ||||||||
| RelBase | 69.4 | 72.0 | ||||||
| SCL | 62.8 | |||||||
| SRAN | ||||||||
| STSN | ||||||||
| WReN |
| Mean | Center | 2x2Grid | 3x3Grid | L-R | U-D | O-IC | O-IG | |
| ALANS | ||||||||
| CPCNet | ||||||||
| CNN-LSTM | ||||||||
| CoPINet | ||||||||
| DRNet | 58.2 | |||||||
| MRNet | ||||||||
| PrAE | 69.6 | 69.5 | ||||||
| PredRNet | ||||||||
| RelBase | 65.7 | 64.9 | 75.6 | |||||
| SCL | 70.8 | 71.8 | ||||||
| SRAN | ||||||||
| STSN | ||||||||
| WReN |
Appendix D Datasheets for datasets
In what follows, we provide the description of the introduced datasets following the Datasheets for Datasets template proposed by Gebru et al. Gebru et al. [2021].